one number is 3 more then twice a second number. their sum is 27 find th enumber

Answer by mananth(16946)

let number be x
other number = 2x+3
sum = 27
...
x+2x+3=27
3x+3=27
-3
3x=24
/3
x=8
..
2x+3 = 16+3 = 19
..
numbers are 8 & 19
m.ananth@hotmail.ca